OpenStudy (anonymous):

Pleas help me Evaluate this Expression?(: -3/6 X -4 + (7 - 8) + 12^2z The -3/6 is a fraction and the X is a multiplication sign if there was any confusion on this question. lol!

OpenStudy (anonymous):

where duh heck did the z came from!?!?

OpenStudy (anonymous):

oh my bad I didn't man to put that.

OpenStudy (anonymous):

lol it's ok

OpenStudy (anonymous):

just use foil i gtg to school now so sry mate lol

OpenStudy (anonymous):

Do you know the BIDMAS mnemonic? Check out - http://www.bbc.co.uk/bitesize/ks3/maths/number/order_operation/revision/2/

OpenStudy (anonymous):

back no not rly lol

OpenStudy (anonymous):

-0.5(-4) -1+144 = 2 + 143 = 145